Calculate the volume if hydrogen under normal conditions, which is released under the influence of excess hydrochloric acid on 22.4 g of the interaction proceeds according to the equation:Fe+2HCl==FeCl2+H2
1
Expert's answer
2020-01-31T04:34:47-0500

We solve by the formula:

n = m / M

where m is the mass of the element

M is the molar mass, it always numerically coincides with the molecular mass (or atomic mass - if the substance does not consist of molecules, but of atoms), we look at the periodic table, 55.847 for iron. So:

n = 22.4 / 55.847 = 0.4 mol of iron

We find n hydrogen by the reaction equation: 1 / 0.4 = 1 / x, x = 0.4 mol

Find the volume of hydrogen:

n = V / Vm = 0.4 * 22.4 = 8.96 l.